10. An electron is separated from a proton by a distance of 1.00 x 10~10 m. What is the magnitude of the electric force between them? 02.30 10-18 N 02.2би 10-10 N 2.30 x 10-8 N 14.2 N
Answer #1

we have the formula

F=((1/(4pi\varepsilono))*(q1q2)/r2

1/4pi\varepsilon=9*109

q1=-q2=-1.6*10-19C

F=9*109*(-1.6*10-19)2/(1*10-10)2=2.30*10-8N

so correct option is 2.30*10-8N.